Solution for: -1 x log(2) = log (x)
Answer by stanbon(75887)

Solution for: -1 x log(2) = log (x)
----
-0.3010 = log(x)
-----
x = 10^(-0.3010)
-----
x = 1/2
=============
Cheers,
Stan H.
=============